# Diff4MMLiTS: Advanced Multimodal Liver Tumor Segmentation via Diffusion-Based Image Synthesis and Alignment

### v1

@@ IV-C Overall Performance
Table III: Results of Diff4MMLiTS Equipped with Different Backbone Models. The Best Results are Highlighted in Bold.
@@
Compared to the U-Net and its variants, our method enhances the DSC by 1.74% over the second-best SwinUNETR. When compared with other multimodal medical image segmentation models, our method enhances DSC by 1.29% aover the second best SegResNet. Moreover, Diff4MMLiTS significantly reduces false positives and false negatives, as reflected in the improved SE and PRE. This indicates enhanced tumor detection capability and a diminished likelihood of missed or incorrect detections. Overall, our proposed method has excellent superiority and achieves the best performance in the multimodal liver tumor segmentation task.
IV-D Generalizability and Adaptability of Diff4MMLiTS
We evaluate the results of our proposed method on publicly available external datasets to verify that the model trained with Diff4MMLiTS can effectively generalize to out-of-distribution data without the need for retraining on the new dataset. All methods are trained on mmLiTs and tested on lesion samples selected from LiTS. The comparison results as shown in Table II. Compared to nnUNet fully trained on real data, Diff4MMLiTS with the synthesis strategy achieves a 16.12% improvement in DSC. This implies that training models with such reliable synthetic images can effectively mitigate the risk of overfitting to in-distribution samples, thereby enhancing their ability to generalize to out-of-distribution samples. This underscores the potential of the proposed method as a promising solution for liver tumor screening.
To further evaluate the adaptability of Diff4MMLiTS, we use three backbone models in the MS module, namely U-Net, AttentionUNet, and nnUNet, with results presented in Table III. The findings indicate that our framework adapts seamlessly to all backbones, achieving notable performance improvements. Compared to segmentation models trained solely on real data, those employing the hybrid training strategy show improvements of 1.95%, 6.58%, and 2.68% in DSC. This identifies the adaptability of our framework to different backbone models and the effectiveness of the hybrid training strategy in enhancing segmentation performance.
Table IV: Results Using Only a Portion of Real Diseased CT and Corresponding Normal CT Training. Note that uni is Short for Unimodal and multi is Short for Multimodal.
@@ IV-D Generalizability and Adaptability of Diff4MMLiTS
IV-E Effectiveness of Multimodal Synthesis Strategy
As illustrated in Table IV, we further evaluate the performance of the synthesis strategy on multimodal and unimodal segmentation methods. In all quantity settings, we employ nnUNet as the segmentation model architecture. The performance of unimodal segmentation models typically rely heavily on the quantity and diversity of training data, but our method requires only a small subset of multimodal data to exceed the performance of models trained on fully annotated unimodal data. The results indicate that when employing merely 10% of the data (real samples and inpaint-based normal multimodal CTs from three patients) as the training set for the segmentation model, it yields a DSC that is 2.63% higher than that of unimodal segmentation models. Furthermore, even when compared to multimodal segmentation models, our method achieves comparable results using only 70% of the paired training data. These findings confirm the effectiveness of our Diff4MMLiTS in enhancing the performance of liver tumor segmentation models.
Table V: Results of Diff4MMLiTS Equipped with Different Inpainting or Synthesis Methods. \$ Denotes the Replacement of the Inpainter and # Denotes the Replacement of the Diffusion Model. The Best Results are Highlighted in Bold.
@@ IV-E Effectiveness of Multimodal Synthesis Strategy
IV-F Contribution of Different Modules to the Framework
In our framework, each stage is trained independently, with the output of one stage serving as the input for the subsequent one. This sequential training method ensures that each component is optimized for its specific role before integration into the overall framework. To comprehensively validate the importance of each module, we explore alternative methods by replacing the inpainter and the diffusion-based synthesis module. Specifically, we replace the inpainter in NCG with a median filter. On the other hand, we replace the proposed latent diffusion-based MCS module with CUT-GAN[35], which constructs the domain of images to be transformed by concatenating normal CTs with randomly generated tumor masks, and learns the mapping between this domain and the domain of the CT image with real tumor.
The impact of these alternatives on the overall framework’s performance is shown in Table V. The results indicate that the framework with the original inpainter and synthesizer achieves the best performance. Compared to the proposed NCG, the CT images generated by the median filter introduce noise and disrupt the original details, thereby affecting the quality of subsequent tumor synthesis. Additionally, the tumor images synthesized by CUT-GAN are overly smooth, making the complementary features between multimodal tumors less pronounced, which, in turn, affects the performance of the segmentation model. In contrast, the inpainter and synthesizer proposed in our original framework not only preserve background details but also generate diverse and high-quality multimodal liver tumors, playing a crucial role in the liver tumor segmentation task.
V Conclusion

### v2

@@ 3.4 Generalizability and Adaptability of Diff4MMLiTS
Table 3: Results of Diff4MMLiTS equipped with different backbone models. The best results are highlighted in bold.
@@
3.5 Effectiveness of Multimodal Synthesis Strategy
As shown in Table 4, we evaluate the effectiveness of our synthesis strategy on both multimodal and unimodal segmentation methods, using nnUNet as the segmentation backbone in all settings. Our proposed method achieves superior performance using only a small subset of multimodal data. Notably, with just 10% of the data (real samples and inpainted multimodal CTs from three patients), our method outperforms the unimodal model trained on fully annotated data by 2.63% in DSC. Furthermore, compared to fully supervised multimodal models summarized in Table 1, our method achieves comparable performance using only 70% of the training data. These results demonstrate the efficiency and effectiveness of Diff4MMLiTS in enhancing liver tumor segmentation.
Table 4: Results using only a portion of real diseased CT and corresponding normal CT training. Note that uni is short for unimodal and multi is short for multimodal.
@@ 3.5 Effectiveness of Multimodal Synthesis Strategy
Table 5: Results of Diff4MMLiTS equipped with different inpainting or synthesis methods. \$ and # denotes the replacement of the inpainter and synthesizer, respectively. The best results are highlighted in bold.
@@
3.6 Contribution of Different Modules to the Framework
To comprehensively validate the importance of each module, we explore alternative designs. Specifically, we replace the inpainter in NCG with a median filter. For MCS, we replace the LDM in MCS with CUT [cutgan], which constructs the source domain by concatenating normal CTs with randomly generated tumor masks and learns a mapping to the target domain of real tumor CTs. The impact of these alternatives on overall performance is shown in Table 5. The results indicate that the framework with the original inpainter and synthesizer achieves the best performance. In comparison, CTs generated by the median filter introduce noise and disrupt the original details, thereby affecting the quality of subsequent tumor synthesis. Additionally, the tumor images synthesized by CUT are overly smooth, making the complementary features between multimodal tumors less pronounced, which, in turn, affects the performance of the segmentation model. In contrast, the inpainter and synthesizer proposed in our original framework not only preserve background details, but also generate reliable multimodal liver tumors, playing a crucial role in the liver tumor segmentation task.
4 Conclusion
@@
5 Acknowledgments
This study was supported by the National Key Research and Development Program of China (2023YFC2415400); the National Natural Science Foundation of China (T2422012, 62071210); the Guangdong Basic and Applied Basic Research (2024B1515020088); the Shenzhen Science and Technology Program (RCYX2021
0609103056042); the High Level of Special Funds (G030230001, G03034K003).